OpenStudy (anonymous):

During a physics experiment, a ball is dropped from a height of 2 m above the ground. Each student in the class has his/her own stopwatch and records the time it takes for the ball to drop that height. The experiment is repeated 10 times and the results are averaged. An average time of 0.70 seconds is calculated from the experimental results. a) Calculate the acceleration due to gravity based on the experimental result.

OpenStudy (asnaseer):

use the equation:\[s=ut+\frac{1}{2}gt^2\]

OpenStudy (asnaseer):

in your case, you know u=0

OpenStudy (asnaseer):

s represents the distance travelled and g is the acceleration due to gravity

OpenStudy (anonymous):

s= 1/2gt^2

OpenStudy (anonymous):

2*2=g*0.7^2 g=4,081